OpenStudy (toxicsugar22):

Calculate the amount of heat required to raise the temperature of a 39g sample of water from 5∘C to 20∘C.

OpenStudy (anonymous):

Q= mc (T2 - T1)

OpenStudy (anonymous):

Just Subs ....

OpenStudy (anonymous):

i have to go...byee

OpenStudy (anonymous):

you wrote from 5 to 20 so (20-5) = difference in temperature..delta T Q = mC(T2-T1) here C is "specific heat capacity of water" = 4.18J/(C*g) Q = 39*4.18*15 ~ 2500J or 2.5 kJ (look carefully, I may made a mistake, but I'm pretty sure)

OpenStudy (anonymous):

to convert from celsius to fahrenheit use this formula: C * (9/5) + 32 = F

OpenStudy (toxicsugar22):

An adult eats food whose nutritional energy totals approximately 2200 Cal per day. The adult burns 2000 Cal per day. If 1 lb of fat is stored by the body for each 14.6×10^3kJ of excess nutritional energy consumed, how long will it take this person to gain 1 lb?